# Universal framework for anisotropic particles with resonance laws and splitting

### v5

@@ Abstract
Nanophotonics enables unprecedented control over light-matter interactions, yet conventional isotropic materials limit the spectral range and richness of mode responses in subwavelength structures. Anisotropic subwavelength particles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions. Here, we establish a universal quasistatic framework describing the eigenmodes and resonance conditions of uniaxial and biaxial nanoparticles. Closed-form solutions reveal axial-permittivity sum rules and a fundamental material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns. Generalizing the theory to ellipsoids provides geometric tunability of the multispectral response, while analytic predictions of quality factors elucidate how anisotropy governs mode localization and energy loss. Full-wave simulations of h-BN and -MoO3 nanoparticles confirm the theory and highlight its predictive power. This framework unifies the understanding of anisotropic nanostructures across optics, thermal transport, and magnetism, opening pathways to a new generation of photonic devices with tunable multispectral response and controlled emission patterns, including directional detectors and spectrally selective mid infrared sensors.
@@ Results
Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and in full-wave electrodynamic simulations. While our theory applies to any uniaxial and biaxial material, we have chosen these materials as they are currently at the forefront of photonics research and enable to achieve high Q factors. We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes and the permittivity model used in Figs. 2,3 [54]. The sphere has a radius of and we situate an oscillating dipole at with a dipole moment of We perform frequency domain simulations with a frequency sweep from (1/s) to (1/s) with 304 frequencies using extensive COMSOL simulations. The dipole moment was chosen to be in the direction of since we assume [67, 72, 31], and prove in the SM, that the mode excitation coefficient depends on so that we need to incorporate only preserving the angular dependence in . This, together with the fact that the dipole has only components, ensures that the degenerate modes will add constructively, as is explained in the Methods Section. Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflect the presence of higher spatial frequencies near the dipole, which excite higher-order modes. For a general incoming field, the mode excitation coefficient is given by where is a window function that equals one in the particle volume, and we utilized a dipole source since a far field excites high-order modes with negligible magnitudes; more details are given in the SM. We then calculate the resonance peak widths analytically (from 1/resonance condition) and from the simulations. To demonstrate the mode-coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the axis. In this case, we do not expect the modes to show in the spectrum since in order to excite them a component of the dipole location is required. In Fig. 4 (top panel) we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of compared to the surface integral of over the uniaxial particle envelope calculated in COMSOL with excellent agreement. Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows) and mode degeneracy for uniaxial spheres. We also obtained the peak widths analytically from and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to is altered due to the peak on the right, and therefore we haven’t analyzed it). Clearly, there is remarkable agreement between them, which showcases our important ability to analytically predict the Q factors given by We then plot for all the observed peaks inside the particle calculated in COMSOL and obtained analytically from the modes (Methods). While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement. Moreover, we plot for a dipole located at the same distance as the previous case and oriented along (31nm). As we predicted the modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation. This highlights the directional nature of the modes as opposed to isotropic particles.
@@ Acknowledgement
E. E. Narimanov, A. Niv, Y. Mazor, J. Jeffet, and Y. Reches are acknowledged for the insightful comments. We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
@@ I.1 Correspondence between the derived modes and the radiation patterns
For the second term we employ the same identity to get
Then, we utilize the spherical-harmonic identity [80]
@@
All in all we get:
We see that when since the overall scaling is of in agreement with the quasistatic analysis. For since and the first term, which is transverse and has scaling as expected, dominates. Thus, one can conclude from the superposition of that participate in the quasistatic anisotropic modes, that the radiation pattern will be of the form:
@@ I.3.1 Theory
Supplementary Material
Completeness of the solution
We now prove that the derived uniaxial-sphere modes are all the eigenmodes. Let us assume that there is an additional eigenstate besides the four modes for each . Since the isotropic-sphere modes span the space inside the sphere, this eigenstate is a superposition of the remaining isotropic-sphere modes
An eigenstate must satisfy anisotropic Laplace’s equation inside the inclusion, which reads:
Since it is a superposition of isotropic-sphere modes we write:
However, all these modes have orders of that are 2 or higher (including from ) and each function is fundamentally different from the others. Hence, such an eigenstate does not exist, and we have a contradiction. Conclusion: the derived uniaxial sphere modes are all the modes. Note that the uniaxial mode basis does not span space as in the standard case.
Similarly, we write for a biaxial particle
If we get which is a uniaxial medium, and we therefore exclude this case. All the modes are isotropic (from the boundary conditions) and the modes satisfy when Even if such modes exist, their axial permittivities must satisfy both Laplace’s equation (in a non-trivial manner) and the boundary condition, which imposes strong constraints. As a result, their potential contribution is typically weak. Therefore, we expect that the dominant biaxial modes satisfy and considered all such possible modes in the main text.
Eigenmode expansion for an anisotropic inclusion
Eigenstate expansions have been derived for an isotropic inclusion and for an inclusion with only one axial permittivity that differs from the host medium permittivity [79, 67]. Here we expand the quasielectrostatic potential for an anisotropic inclusion. For concreteness, we consider a uniaxial inclusion and the derived eigenstates. Generalizing it to a biaxial inclusion readily follows.
We start by writing Poisson’s equation for a medium with an anisotropic inclusion:
where is a step function that equals 1 inside the inclusion, is the inclusion permittivity tensor, and is the host medium permittivity. Now we move one term to the lhs in order to be able to define an eigenvalue equation.
where . The Green function in this case will be for a medium with an inclusion with anisotropy in the axis and axial permittivities that are equal to the host medium permittivity in the other axes.
where we defined to adjust it to an eigenvalue equation with an operator operating on and we performed integration by parts.
We express the electric potential and expand it using the anisotropic eigenfunctions defined by Laplace’s equation. In the following is defined for the particle with anisotropy in the axis. We write:
Now we expand using the eigenstates of a particle with anisotropy in the axis:
where is the electric potential for the dipole in free space. All in all we get:

### v6

@@ Abstract
Nanophotonics enables unprecedented control over light-matter interactions, yet conventional isotropic materials limit the spectral range and richness of mode responses in subwavelength structures. Anisotropic subwavelength particles—ubiquitous in natural and engineered systems—offer new degrees of freedom that couple geometry and material properties, unlocking previously inaccessible spectral regions. Here, we establish a universal full-wave framework describing the eigenmodes and resonance conditions of uniaxial and biaxial nanoparticles. Closed-form solutions reveal axial permittivity sum rules and a fundamental material-anisotropy-induced symmetry breaking, manifesting as resonance splitting and novel radiation patterns. Generalizing the theory to ellipsoids provides geometric tunability of the multispectral response, while analytic predictions of quality factors elucidate how anisotropy governs mode localization and energy loss. Full-wave simulations of h-BN and -MoO3 nanoparticles, together with our recently reported experimental observations, confirm the theory and highlight its predictive power. This framework unifies the understanding of anisotropic nanostructures across optics, thermal transport, and magnetism, opening pathways to a new generation of photonic devices with tunable multispectral response and controlled emission patterns, including directional detectors and spectrally selective mid infrared sensors.
@@ Results
Finally, we demonstrate our theory for spheres and ellipsoids composed of hBN and in full-wave electrodynamic simulations. While our theory applies to any uniaxial and biaxial material, we have chosen these materials as they are currently at the forefront of photonics research and enable to achieve high Q factors. We consider a subwavelength uniaxial hBN sphere with the equal permittivity axes and the permittivity model used in Figs. 2,3 [54]. The sphere has a radius of and we situate an oscillating dipole at with a dipole moment of We perform frequency domain simulations with a frequency sweep from (1/s) to (1/s) with 304 frequencies using extensive COMSOL simulations. The dipole moment was chosen to be in the direction of since we assume [67, 72, 31], and prove in the SM, that the mode excitation coefficient depends on so that we need to incorporate only preserving the angular dependence in . This, together with the fact that the dipole has only components, ensures that the degenerate modes will add constructively, as is explained in the Methods Section. Note that the coefficient’s dependence on the mode fields at the dipole location, with the rapid decay of the high-order modes away from the particle, reflect the presence of higher spatial frequencies near the dipole, which excite higher-order modes. For a general incoming field, the mode excitation coefficient is given by where is a window function that equals one in the particle volume, and we utilized a dipole source since a far field excites high-order modes with negligible magnitudes; more details are given in the SM. We then calculate the resonance peak widths analytically (from 1/resonance condition) and from the simulations. To demonstrate the mode-coefficient dependency on the dipole location and orientation, we also perform the same calculation for a dipole oriented along the axis. In this case, we do not expect the modes to show in the spectrum since in order to excite them a component of the dipole location is required. In Fig. 4 (top panel) we present the analytically-predicted peaks in the scattering spectrum obtained from the maxima of compared to the surface integral of over the uniaxial particle envelope calculated in COMSOL with excellent agreement. Importantly, these simulation results for the scattering spectrum confirm our predictions of the resonance splitting (marked by arrows) and mode degeneracy for uniaxial spheres. We also obtained the peak widths analytically from and compared them to the ones calculated from the simulations (as can be seen, the width of the peak, which corresponds to is altered due to the peak on the right, and therefore we haven’t analyzed it). Clearly, there is remarkable agreement between them, which showcases our important ability to analytically predict the Q factors given by We then plot for all the observed peaks inside the particle calculated in COMSOL and obtained analytically from the modes (Methods). While the COMSOL fields include the incoming field from the oscillating dipole, close to a resonance this is typically negligible and indeed we observe very good agreement. Moreover, we plot for a dipole located at the same distance as the previous case and oriented along (31nm). As we predicted the modes disappeared from the spectrum, verifying the dependency of the mode excitation coefficient on the dipole location and orientation. This highlights the directional nature of the modes as opposed to isotropic particles. In Fig. 5 we compared the radiation pattern from the full-wave closed-form solution corresponding to the mode with the resonance condition outside a uniaxial particle to a full-wave COMSOL simulation of the scattered field of a uniaxial particle with nm and excited by a current source similarly to Fig. 4 (top) in the radiation zone in the radiation zone . Note that the COMSOL simulations utilized the axial symmetry and are therefore presented on half a sphere. As can be seen, the radiation-pattern results show very good agreement.
@@ Acknowledgement
E. E. Narimanov, A. Niv, Y. Mazor, J. Jeffet, Y. Reches, and I. Kalyan are acknowledged for the insightful comments. We acknowledge the support from the ISF grant 2312/20 and a grant from The Ministry of Aliyah and Integration.
@@ I.1 Correspondence between the derived modes and the radiation patterns
For the second term we arrive at
where we have used the vector identity for Then, we utilize the spherical-harmonic identity [80]
@@
All in all, we get:
We see that when since the overall scaling is of in agreement with the quasistatic analysis. For since and the first term, which is transverse and has scaling as expected, dominates. Thus, one can conclude from the superposition of that participate in the quasistatic anisotropic modes, that the radiation patterns are of the form: